# STELAR-VISION: Self-Topology-Aware Efficient Learning for Aligned Reasoning in Vision

### v2

@@ Abstract
Vision-language models (VLMs) have made significant strides in reasoning, yet they often struggle with complex multimodal tasks and tend to generate overly verbose outputs. A key limitation is their reliance on chain-of-thought (CoT) reasoning, despite many tasks benefiting from alternative topologies like trees or graphs. To address this, we introduce STELAR-Vision, a training framework for topology-aware reasoning. At its core is TopoAug, a synthetic data pipeline that enriches training with diverse topological structures. Using supervised fine-tuning and reinforcement learning, we post-train Qwen2VL models with both accuracy and efficiency in mind. Additionally, we propose Frugal Learning, which reduces output length with minimal accuracy loss. On MATH-V and VLM_S2H, STELAR-Vision improves accuracy by 9.7% over its base model and surpasses the larger Qwen2VL-72B-Instruct by 7.3%. On five out-of-distribution benchmarks, it outperforms Phi-4-Multimodal-Instruct by up to 28.4% and LLaMA-3.2-11B-Vision-Instruct by up to 13.2%, demonstrating strong generalization. Compared to Chain-Only training, our approach achieves 4.3% higher overall accuracy on in-distribution datasets and consistently outperforms across all OOD benchmarks.
Datasetshttps://huggingface.co/collections/Stellar-Neuron/stelar-vision-collection
Project Pagehttps://stellar-neuron.github.io/stelar-vision/
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved reasoning capabilities, with models like GPT-o3 achieving strong performance on complex mathematical and scientific tasks. This progress has extended into the multimodal domain through vision-language models (VLMs) such as GPT-4o (Hurst et al. 2024), GPT-4o-mini (OpenAI 2024), and Qwen2.5-VL (Bai et al. 2025). Despite the recent advances, there is still room of improvement in open-sourced VLMs when tackling complex vision-based reasoning tasks (e.g., math and science questions), and the path to enhance their abilities under an affordable training budget remains under-explored.
@@
To address this, we begin by analyzing VLMs’ reasoning behaviors and find that the popular models, both open-source and closed-source, tend to default to the chain-of-thought (CoT) (Wei et al. 2023) generation. This behavior reflects the prevailing trend in their training data, which is overwhelmingly dominated by CoT-style reasoning samples. However, our empirical analysis reveals that different questions benefit from different reasoning topologies, such as Chain, Tree, or Graph structures (Figure 1). The benefits of diverse reasoning topologies have yet been well studied or effectively incorporated into existing training pipelines. Moreover, even state-of-the-art reasoning VLMs tend to generate unnecessarily verbose responses, i.e., “overthinking”, which increases the computational cost and makes real-time applications less viable. We find that there is a correlation between the topological reasoning structures and the output sequence length, thus providing an insight to the overthinking problem created by the CoT reasoning.
@@
We introduce Self-Topology-Aware-Efficient-Learning for Aligned Reasoning in Vision, STELAR-Vision, a training framework for topology-aware vision-language reasoning. The central to this framework is TopoAug, a synthetic data generation pipeline that produces augmented reasoning topologies with diverse structures, including $\{Chain,Tree,Graph\}$. Specifically, we generate a set of question-answer responses, where each question is repeatedly answered by different topological reasoning structures. For each question, we assign a preferred topology through an automated annotation process. The generated responses and their labels are used to post-train VLMs via supervised fine-tuning (SFT) and Reinforcement Learning (RL) (Meng, Xia, and Chen 2024). Regarding the learning process, we observe that RL further amplifies the performance gains introduced by topological augmentation. Moreover, we propose to further increase efficiency with Frugal Learning by promoting shorter responses. We show that only by leveraging the benefits of augmented reasoning topologies, can the model generate concise responses and maintain high accuracy with minimal performance drops, while the counterpart model trained only on CoT-style data fails to gain efficiency without incurring a greater loss in accuracy.
@@
The augmented topologies expand the exploration space, allowing RL to discover higher-quality optima and leading to superior performance improvement. In contrast, the model trained with CoT-style data only sees diminishing returns. We also test our models on five out-of-distribution datasets, and it consistently outperform its base model, while consuming fewer tokens. which suggests strong generalization of our proposed STELAR-Vision.
@@ 2.1 Topological Reasoning in Language and Vision Models
These methods, however, often rely on rule-based topology generations through prompt engineering or sampling, and are limited to language-only settings. In contrast, our framework automatically generates diverse topological structures and trains a VLM to adaptively select the optimal one per instance during decoding, enabling more flexible and generalizable reasoning.
@@ 2.2 Reinforcement Learning for LLM and VLM Reasoning
Reward-based methods such as PPO (Schulman et al. 2017), RPO (Yin et al. 2024), and GRPO (Shao et al. 2024) rely on explicit rewards, while reward-free approaches like DPO (Rafailov et al. 2024), SimPO (Meng, Xia, and Chen 2024), and ORPO (Hong, Lee, and Thorne 2024) achieve comparable results without reward modeling. These methods are widely applied to mathematical reasoning, long-horizon tasks, and instruction tuning.
@@ 2.3 Curriculum Learning and Structured Reasoning
Curriculum learning (Bengio et al. 2009) trains models on progressively harder tasks, aiding structured reasoning. Recent works extend this idea: Xi et al. (Xi et al. 2024) use reverse curricula for LLMs, Zhao et al. (Zhao et al. 2024) propose Auto-CEI for ability-aligned sampling, and Ma et al. (Ma, Jiang, and Huang 2025) design logic-guided curricula for in-context learning. Fine-tune-CoT (Ho, Schmid, and Yun 2023) learns CoT from larger reasoning models.
@@ 3.1 Constructing Topology-Aware Responses
We use two math datasets: MATH-V (Wang et al. 2024a) (3,040 visual problems) and VLM_S2H (Park et al. 2025) (7,000 logic puzzles), each pairing images with questions, giving diverse samples requiring high-level reasoning
TopoAug
We generate Topology-Aware responses using topologies $T=\{Chain,Tree,Graph\}$, prompted via Qwen2-VL-7B-Instruct (Wang et al. 2024b) and GPT-4o-Mini (OpenAI 2024a) with extensive degrees of freedom in maximum depth, number of children, and number of neighbors. Please see the supplementary for detailed prompts.
Topology and Outcome Labels
Each response $r$ has an Outcome Label $\mathcal{H}_{r}\in\{0,1\}$, and is assigned label 1 if correct and 0 otherwise. Each question-topology pair gets a Topology Label $\mathcal{F}_{q,t}=\frac{N_{\text{correct}}(q,t)}{N_{\text{total}}(q,t)}$ based on accuracy, where $N_{\text{correct}}(q,t)$ is the number of correct responses using topology $t$ for question $q$, and $N_{\text{total}}(q,t)$ is the total number of responses generated using $t$.
Problem Difficulty Segmentation
Problems are labeled Easy (>>85th percentile), Hard (<<15th), or Medium based on topology score distributions.
@@ 3.2 Analysis: Topological Reasoning Structures
where $N_{Q}$ is the total number of questions, and $\mathds{1}$ is the indicator function. For each question, the topology with the highest Topology Label $\mathcal{F}_{q,t}$ wins. We first analyze which topological reasoning structure works the best for the questions. The Win Rate statistics is presented in Table 1.
@@
Figure 3: Comparison of topology accuracy across subjects: Accuracy of ChainChain, TreeTree, and GraphGraph reasoning topological structures per subject of MATH-V dataset. ChainChain remains the best overall reasoning structure, while TreeTree, and GraphGraph perform better in at reasoning subjects such as “graph theory” or “statistics”.
@@ 3.3 STELAR-Vision Post-Training
We follow prior work (Chu et al. 2025; Zhai et al. 2024) by initializing reinforcement learning (RL) from an SFT checkpoint, which improves both in-distribution accuracy and OOD generalization. We adopt SimPO (Meng, Xia, and Chen 2024) for its simplicity and alignment with test-time behavior. Its objective encourages preferred responses over less-preferred ones, and is defined as
@@
Llama-3.2-11B-Vision-Instruct (Meta 2024) 22.0 10.0 18.0 35.0 37.8 22.2 10.7 24.8
@@
Variant 1: STELAR-Vision-Short. We first introduce a filter that identifies “short and correct” responses as preferred targets, where the outcome Label $\mathcal{H}_{r}=1$ and the generated token length falls below a 25% percentile threshold. The model undergoes training with both supervised fine-tuning (SFT) and reinforcement learning (RL). During RL training, we designate “short and correct” responses as winners and incorrect responses as losers in the preference optimization process. We also train a model with Chain-Only data with the same process and we denote it as Chain-Only-Short.
Variant 2: STELAR-Vision-Short. A natural alternative approach to encourage shorter and more efficient responses involves explicitly penalizing lengthy outputs. Building upon STELAR-Vision-Short, we investigate an alternative configuration, STELAR-Vision-Short, which maintains the preference for “short and correct” responses while treating both “incorrect responses” and “correct yet lengthy” responses as losers during RL preference training. This design encourages the model to avoid excessive token generation while maintaining accuracy in reaching the correct final answer.
@@ 4.1 Experimental Setup
We use In-Distribution (ID) datasets from the Method Section 3, splitting them into train/test set. We include 85K training samples from MATH-V and 160K samples from VLM_S2H. OKVQA and A-OKVQA provide 18k and 20K VQA pairs for training, with A-OKVQA adding rationales to training. We also use 17k instruction-following examples from LLaVA-150k for multimodal training. performed on the test sets of MATH-V and VLM_S2H.
For Out-of-Distribution (OOD) evaluation, we use five recent benchmarks with three math oriented datasets, Geometry3K (Lu et al. 2021), We-Math (Qiao et al. 2024) and PolyMath (Gupta et al. 2024), a STEM dataset SciBench (Wang et al. 2024c) and a generic logic reasoning dataset LogicVista (Xiao et al. 2024). In Table 3, we show the evaluation datasets as well as their respective sizes. Collectively, we gather 2,425 evaluation samples.
@@
Dataset Subject Question Type Sample Size
@@
Table 3: Summary of Evaluation Datasets. VLM_S2H and MATH-V are in-distribution datasets, while others are out-of-distribution
@@ 4.2 Overall Evaluation Results
We use STELAR-Vision to denote models trained with both phases of post-training, the -SFT suffix indicates models trained with supervised fine-tuning only, and -RL-ONLY indicates reinforcement learning directly from the base model without SFT. Results are averaged across 3 random seeds.
@@
Results on out-of-distribution (OOD) benchmarks highlight the strong generalization of STELAR-Vision. It outperforms its base model Qwen2VL-7B-Instruct on all five OOD datasets and trails GPT-4o by only 1.2% on PolyMath. Compared to recently popular model Phi-4-Multimodal-instruct, STELAR-Vision achieves significantly higher accuracy across all OOD tasks, including a +28.4% gain on the spatial reasoning benchmark Geometry3K. It also surpasses other strong open-source models such as LLaVA-v1.6, LLaMA-3.2-11B-Vision, MiniCPMv2.6-8B, and InternVL3-9B on most benchmarks, demonstrating the impact of our TopoAug-enhanced training. STELAR-Vision also outperforms baseline on LogicVista, which includes non-math tasks like 26.4% diagram and 18% spatial.
@@ 4.3 Ablation Studies
These findings collectively demonstrate that STELAR-Vision benefits not only from data distillation or pattern memorization but also from a genuine ability to adaptively select optimal reasoning topologies based on the task.
Model VLM_S2H MATH-V Overall
@@
4.4 Efficiency Gains from Frugal Learning
As shown in Table 5, STELAR-Vision-Short reduces token length by 101 (ID) and 24.5 (OOD), while still outperforming Qwen2VL-7B-Instruct by 2.5%. However, it shows a 2.9% accuracy drop compared to full STELAR-Vision. In contrast, STELAR-Vision-Short yields inconsistent length reduction and greater performance loss, likely due to conflicting optimization signals that penalize correct outputs.
Chain-Only-Short also shortens outputs but suffers from even larger accuracy drop. Notably, Chain-Only models fine-tuned with RL often generate overly verbose responses, even with Frugal Learning—highlighting the advantage of TopoAug in promoting concise and effective reasoning.

### v3

@@ Abstract
Vision-language models (VLMs) have made significant strides in reasoning, yet they often struggle with complex multimodal tasks and tend to generate overly verbose outputs. A key limitation is their reliance on chain-of-thought (CoT) reasoning, despite many tasks benefiting from alternative topologies like trees or graphs. To address this, we introduce STELAR-Vision, a training framework for topology-aware reasoning. At its core is TopoAug, a synthetic data pipeline that enriches training with diverse topological structures. Using supervised fine-tuning and reinforcement learning, we post-train Qwen2VL models with both accuracy and efficiency in mind. Additionally, we propose Frugal Learning, which reduces output length with minimal accuracy loss. On MATH-V and VLM_S2H, STELAR-Vision improves accuracy by 9.7% over its base model and surpasses the larger Qwen2VL-72B-Instruct by 7.3%. On five out-of-distribution benchmarks, it outperforms Phi-4-Multimodal-Instruct by up to 28.4% and LLaMA-3.2-11B-Vision-Instruct by up to 13.2%, demonstrating strong generalization. Compared to Chain-Only training, our approach achieves 4.3% higher overall accuracy on in-distribution datasets and consistently outperforms across all OOD benchmarks. We’ve released datasets, and code will be available.
footnotetext: Preprint
Project and Datasets: https://stellar-neuron.github.io/stelar-vision/
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved reasoning capabilities, with models like GPT-o3 achieving strong performance on complex mathematical and scientific tasks. This progress has extended into the multimodal domain through vision-language models (VLMs) such as GPT-4o (OpenAI et al. 2024), GPT-4o-mini (OpenAI 2024b), and Qwen2.5-VL (Bai et al. 2025). Despite the recent advances, there is still room of improvement in open-sourced VLMs when tackling complex vision-based reasoning tasks (e.g., math and science questions), and the path to enhance their abilities under an affordable training budget remains under-explored.
@@
To address this, we begin by analyzing VLMs’ reasoning behaviors and find that the popular models, both open-source and closed-source, tend to default to the chain-of-thought (CoT) (Wei et al. 2023) generation. This behavior reflects the prevailing trend in their training data, which is overwhelmingly dominated by CoT-style reasoning samples. However, our empirical analysis reveals that different questions benefit from different reasoning topologies, such as Chain, Tree, or Graph structures (Please see Figure 1). The benefits of diverse reasoning topologies have yet been well studied or effectively incorporated into existing training pipelines. Moreover, even state-of-the-art reasoning VLMs tend to generate unnecessarily verbose responses, i.e., “overthinking", which increases the computational cost and makes real-time applications less viable. We find that there is a correlation between the topological reasoning structures and the output sequence length, thus providing an insight to the overthinking problem created by the CoT reasoning.
@@
We introduce Self-Topology-Aware-Efficient-Learning for Aligned Reasoning in Vision, STELAR-Vision, a training framework for topology-aware vision-language reasoning. The central to this framework is TopoAug, a synthetic data generation pipeline that produces augmented reasoning topologies with diverse structures, including $\{Chain,Tree,Graph\}$. Specifically, we generate a set of question-answer responses, where each question is repeatedly answered by different topological reasoning structures. For each question, we assign a preferred topology through an automated annotation process. The generated responses and their labels are used to post-train VLMs via supervised fine-tuning (SFT) and Reinforcement Learning (RL) (Meng et al. 2024). Regarding the learning process, we observe that RL further amplifies the performance gains introduced by topological augmentation. Moreover, we propose to further increase efficiency with Frugal Learning by promoting shorter responses. We show that only by leveraging the benefits of augmented reasoning topologies, can the model generate concise responses and maintain high accuracy with minimal performance drops, while the counterpart model trained only on CoT-style data fails to gain efficiency without incurring a greater loss in accuracy.
@@
The augmented topologies expand the exploration space, allowing RL to discover higher-quality optima and leading to superior performance improvement. In contrast, the model trained with CoT-style data only sees diminishing returnsWe also test our models on five out-of-distribution datasets and consistently outperform its base model. which suggests strong generalization of our proposed STELAR-Vision.
@@ 2.1 Topological Reasoning in Language and Vision Models
These methods, however, often rely on rule-based topology generations through sampling and are limited to language-only settings. In contrast, our framework automatically generates diverse topological structures and trains a VLM to adaptively select the optimal one per instance during decoding, enabling more flexible and generalizable reasoning.
@@ 2.2 Reinforcement Learning for LLM and VLM Reasoning
Reward-based methods such as PPO (Schulman et al. 2017), RPO (Yin et al. 2024), and GRPO (Shao et al. 2024) rely on explicit rewards, while reward-free approaches like DPO (Rafailov et al. 2024), SimPO (Meng et al. 2024), and ORPO (Hong et al. 2024) achieve comparable results without reward modeling. These methods are widely applied to mathematical reasoning, long-horizon tasks, and instruction tuning.
@@ 2.3 Curriculum Learning and Structured Reasoning
Curriculum learning (Bengio et al. 2009) trains models on progressively harder tasks, aiding structured reasoning. Recent works extend this idea: Xi et al. (Xi et al. 2024) use reverse curricula for LLMs, Zhao et al. (Zhao et al. 2024) propose Auto-CEI for ability-aligned sampling, and Ma et al. (Ma et al. 2025) design logic-guided curricula for in-context learning.
@@ 3.1 Constructing Topology-Aware Responses
We begin with two mathematical datasets MathVision (MATH-V) (Wang et al. 2024a), and VLM_S2H (Park et al. 2025). They provide diverse samples requiring high-level reasoning, suitable to our study. Every sample in each dataset is composed of an image and a text-based problem description. MATH-V contains 3,040 math problems with visual contexts from real-world competitions. VLM_S2H (Park et al. 2025) offers 7,000 structured reasoning puzzles testing logical relations like Progression and Exclusive OR.
TopoAug: Generating Topology-Aware Response
For each question, we prompt these models to repeatedly generate multiple responses across three reasoning topologies, $T=\{Chain,Tree,Graph\}$ with extensive degrees of freedom in maximum depth, number of children, and number of neighbors. Please see the supplementary for detailed prompts.
To ensure diversity in the reasoning topologies and to maintain a balanced distribution of positive and negative samples in our dataset, we use two models with different scale and abilities to fulfill the generation: open-sourced Qwen2-VL-7B-Instruct (Wang et al. 2024b), and commercial GPT-4o-Mini (OpenAI 2024a) for data generation.
Each question gets multiple generated responses via different topology prompting.
Topology Label and Outcome Label
We assign each question three Topology Label $\mathcal{F}_{q,t}$ and assign each response an Outcome Label $\mathcal{H}_{r}$, which are defined as below.
  • Topology Label $\mathcal{F}_{q,t}$: A continuous value in the range $[0,1]$, representing the accuracy that a given topology produces the correct answers for a question. This question-specific label reflects how effectively each reasoning topology solves a given problem. For each problem $q$, we compute the accuracy of the responses from each topology type and assign it as: $\mathcal{F}_{q,t}=\frac{N_{\text{correct}}(q,t)}{N_{\text{total}}(q,t)}$, where $N_{\text{correct}}(q,t)$ is the number of correct responses using topology $t$ for question $q$, and $N_{\text{total}}(q,t)$ is the total number of responses generated using $t$.
  • Outcome Label $\mathcal{H}_{r}$: For each generated response, a binary value $\{0,1\}$ indicating whether it is correct. Each response $r$ is assigned label 1 if correct and 0 otherwise.
Problems Difficulty Segmentation
TopoAug enables us to quantify problem to three difficulty levels based on the distribution of topology labels:
  • Easy: Problems where all three topology scores exceed a quantile threshold (85%).
  • Hard: Problems where all three topology scores fall below a specified quantile threshold (15%) in their respective distributions.
  • Medium: Problems that do not meet the criteria for either hard or easy categories.
@@ 3.2 Analysis: Topological Reasoning Structures
where $N_{Q}$ is the total number of questions, and $\mathds{1}$ is the indicator function. For each question, the topology with the highest Topology Label $\mathcal{F}_{q,t}$ is assigned a win. We first analyze which topological reasoning structure works the best for the questions. The Win Rate statistics is presented in Table 1.
@@
Figure 3: Comparison of topology accuracy across subjects: Accuracy of ChainChain, TreeTree, and GraphGraph reasoning topological structures per subject of MATH-V dataset. ChainChain remains the best overall reasoning structure, while TreeTree, and GraphGraph perform better in at reasoning subjects such as “graph theory" or “statistics".
@@ 3.3 STELAR-Vision Post-Training
We follow prior work (Chu et al. 2025; Zhai et al. 2024) by initializing reinforcement learning (RL) from an SFT checkpoint, which improves both in-distribution accuracy and OOD generalization. We adopt SimPO (Meng et al. 2024) for its simplicity and alignment with test-time behavior. Its objective encourages preferred responses over less-preferred ones, and is defined as
@@
Llama-3.2-11B-Vision-Instruct (Grattafiori et al. 2024; Meta 2024) 22.0 10.0 18.0 35.0 37.8 22.2 10.7 24.8
@@
Variant 1: STELAR-Vision-Short. We first introduce a filter that identifies “short and correct" responses as preferred targets, where the outcome Label $\mathcal{H}_{r}=1$ and the generated token length falls below a 25% percentile threshold. The model undergoes training with both supervised fine-tuning (SFT) and reinforcement learning (RL). During RL training, we designate “short and correct" responses as winners and incorrect responses as losers in the preference optimization process. We also train a model with Chain-Only data with the same process and we denote it as Chain-Only-Short.
Variant 2: STELAR-Vision-Short. A natural alternative approach to encourage shorter and more efficient responses involves explicitly penalizing lengthy outputs. Building upon STELAR-Vision-Short, we investigate an alternative configuration, STELAR-Vision-Short, which maintains the preference for “short and correct" responses while treating both “incorrect responses" and “correct yet lengthy" responses as losers during RL preference training. This design encourages the model to avoid excessive token generation while maintaining accuracy in reaching the correct final answer.
@@ 4.1 Experimental Setup
We use In-Distribution (ID) datasets from the Method Section 3, splitting them into train/test sets with  50K–60K samples for training. OKVQA and A-OKVQA provide 18k and 36k VQA pairs, with A-OKVQA adding rationales and new knowledge types. We also use 17k instruction-following examples from LLaVA-150k for multimodal training. The number of the training samples in Table 3.
Dataset MATH-V VLM_S2H OKVQA A-OKVQA LLava150k-inst
Sample 85K 160K 18K 20K 17K
Table 3: Total sample sizes of datasets used for training.
Evaluation is performed on the test sets of MATH-V and VLM_S2H. For Out-of-Distribution (OOD) evaluation, we use five recent benchmarks with three math oriented datasets, Geometry3K (Lu et al. 2021), We-Math (Qiao et al. 2024) and PolyMath (Gupta et al. 2024), a STEM dataset SciBench (Wang et al. 2024c) and a generic logic reasoning dataset LogicVista (Xiao et al. 2024). In Table 4, we show the evaluation datasets as well as their respective sizes. Collectively, we gather 2,425 evaluation samples.
@@
Dataset Subject Is OOD? Question Type Sample Size
@@
Table 4: Summary of Evaluation Datasets
@@ 4.2 Overall Evaluation Results
We use STELAR-Vision to denote models trained with both phases of post-training, the -SFT suffix indicates models trained with supervised fine-tuning only, and -RL-ONLY indicates reinforcement learning directly from the base model without SFT.
@@
Results on out-of-distribution (OOD) benchmarks highlight the strong generalization of STELAR-Vision. It outperforms its base model Qwen2VL-7B-Instruct on all five OOD datasets and trails GPT-4o by only 1.2% on We-Math. Compared to recently popular model Phi-4-Multimodal-instruct, STELAR-Vision achieves significantly higher accuracy across all OOD tasks, including a +28.4% gain on the spatial reasoning benchmark Geometry3K. It also surpasses other strong open-source models such as LLaVA-v1.6, LLaMA-3.2-11B-Vision, MiniCPMv2.6-8B, and InternVL3-9B on most benchmarks, demonstrating the impact of our TopoAug-enhanced training.
@@ 4.3 Ablation Studies
These findings collectively demonstrate that STELAR-Vision benefits not only from data distillation or pattern memorization but also from a genuine ability to internalize and adaptively select optimal reasoning topologies based on the task.
Model w/ SFT w/ RL VLM_S2H MATH-V Overall
@@
4.4 Efficiency Gains from Frugal Learning
As shown in Table 6, STELAR-Vision-Short reduces token length by 101 (ID) and 24.5 (OOD), while still outperforming Qwen2VL-7B-Instruct by 2.5%. However, it shows a 2.9% accuracy drop compared to the full STELAR-Vision. In contrast, STELAR-Vision-Short yields inconsistent length reduction and greater performance loss, likely due to conflicting optimization signals that penalize correct outputs.
Chain-Only-Short also shortens outputs but suffers from even larger accuracy degradation. Notably, Chain-Only models fine-tuned with RL often generate overly verbose responses, even with Frugal Learning—highlighting the advantage of TopoAug in promoting concise and effective reasoning.